OpenStudy (anonymous):

we can demonstrate that a statement such as (a + b)^2 = a^2 + b^2 is false by saying

hero (hero):

Let a = 1 and b = 2 (1 + 2)^2 = 1^2 + 2^2 3^2 = 1 + 4 9 = 5 False

OpenStudy (anonymous):

thanks dude i suck at math